# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

### v1

@@ Appendix C Additional results
C.1 Detailed results for Section 3.1
@@
We report additional results for Sec. 3.1, showing the AUC for each downstream model (XGBoost, Random forest, Logistic regression, Decision tree). As we can see, the conclusion that curation helps improve downstream performance holds for each of these various downstream models, as is indicated by the green arrows in Tables 6, 7, 8, 9.
@@
cutract (n=20) 74.16 73.34 73.45 74.24 normal-$\uparrow$ 72.22 73.51 $\uparrow$ 66.24 70.45 $\uparrow$ 69.73 70.49 $\uparrow$ 52.61 68.11 $\uparrow$ 66.63 72.39 $\uparrow$ 71.42 72.34 $\uparrow$ 71.18 71.89 $\uparrow$
maggic (n=20) 71.28 59.18 62.92 63.97 normal-$\uparrow$ 60.10 61.25 $\uparrow$ 53.21 55.14 $\uparrow$ 55.61 56.50 $\uparrow$ 50.92 57.23 $\uparrow$ 55.91 59.07 $\uparrow$ 58.20 58.55 $\uparrow$ 55.43 57.71 $\uparrow$
seer (n=20) 90.09 85.22 86.30 86.82 normal-$\uparrow$ 85.86 85.73 78.11 80.52 $\uparrow$ 82.61 82.74 $\uparrow$ 47.57 75.70 $\uparrow$ 77.35 81.45 $\uparrow$ 82.09 83.00 $\uparrow$ 77.63 81.16 $\uparrow$
@@
covid (n=40) 76.71 75.21 75.63 75.70 normal-$\uparrow$ 72.27 73.59 $\uparrow$ 66.64 70.75 $\uparrow$ 75.56 75.39 56.91 70.88 $\uparrow$ 70.47 73.15 $\uparrow$ 73.59 73.63 $\uparrow$ 64.27 69.71 $\uparrow$
cutract (n=40) 74.96 72.31 71.67 73.20 normal-$\uparrow$ 69.92 71.01 $\uparrow$ 64.67 69.21 $\uparrow$ 69.16 69.67 $\uparrow$ 53.41 68.93 $\uparrow$ 61.23 69.21 $\uparrow$ 70.82 71.21 $\uparrow$ 59.99 67.56 $\uparrow$
maggic (n=40) 71.32 60.91 61.80 63.11 normal-$\uparrow$ 59.38 61.50 $\uparrow$ 56.49 58.17 $\uparrow$ 56.50 58.21 $\uparrow$ 48.43 57.84 $\uparrow$ 55.47 59.94 $\uparrow$ 59.82 60.11 $\uparrow$ 56.76 58.92 $\uparrow$
seer (n=40) 90.06 86.96 86.01 86.29 $\uparrow$ 86.73 87.09 normal-$\uparrow$ 83.46 86.62 $\uparrow$ 85.86 85.19 54.71 83.22 $\uparrow$ 83.23 86.81 $\uparrow$ 83.76 84.82 $\uparrow$ 80.34 85.82 $\uparrow$
compas (n=40) 66.66 62.73 68.06 68.13 normal-$\uparrow$ 61.17 62.27 $\uparrow$ 56.05 60.92 $\uparrow$ 59.76 61.03 $\uparrow$ 57.82 64.25 $\uparrow$ 58.89 63.79 $\uparrow$ 61.20 61.25 $\uparrow$ 55.89 60.54 $\uparrow$
adult (n=40) 86.04 83.61 50.26 75.55 $\uparrow$ 46.68 76.63 $\uparrow$ 75.13 81.56 $\uparrow$ 76.10 80.40 $\uparrow$ 68.04 81.08 $\uparrow$ 73.29 81.46 $\uparrow$ 80.35 81.10 $\uparrow$ 75.81 82.11 $\uparrow$
drug (n=40) 82.76 78.81 78.35 79.54 normal-$\uparrow$ 77.33 77.94 $\uparrow$ 74.68 76.64 $\uparrow$ 74.50 77.18 $\uparrow$ 70.25 $7$uparrow 65.86 74.55 $\uparrow$ 74.84 75.26 $\uparrow$ 71.29 75.98 $\uparrow$
covid (n=100) 76.21 75.78 75.86 76.33 normal-$\uparrow$ 73.02 74.40 $\uparrow$ 72.00 74.74 $\uparrow$ 74.64 75.74 $\uparrow$ 65.76 74.69 $\uparrow$ 67.05 75.47 $\uparrow$ 74.00 74.09 $\uparrow$ 72.11 74.44 $\uparrow$
cutract (n=100) 74.59 73.93 72.92 74.73 $\uparrow$ 72.51 73.93 $\uparrow$ 70.18 74.26 $\uparrow$ 71.71 73.69 $\uparrow$ 55.48 70.79 $\uparrow$ 69.79 74.75 normal-$\uparrow$ 71.18 72.47 $\uparrow$ 68.47 73.78 $\uparrow$
maggic (n=100) 71.28 63.06 60.99 63.36 normal-$\uparrow$ 57.97 60.86 $\uparrow$ 58.98 60.76 $\uparrow$ 58.88 60.13 $\uparrow$ 49.66 59.57 $\uparrow$ 57.82 62.24 $\uparrow$ 61.55 61.90 $\uparrow$ 57.46 60.55 $\uparrow$
seer (n=100) 89.94 87.53 86.33 87.31 $\uparrow$ 86.40 87.06 $\uparrow$ 84.59 87.27 $\uparrow$ 85.89 87.00 $\uparrow$ 70.32 85.97 $\uparrow$ 84.30 87.52 $\uparrow$ 85.12 85.82 $\uparrow$ 81.92 86.41 $\uparrow$
compas (n=100) 66.38 63.18 68.44 68.67 normal-$\uparrow$ 59.41 62.00 $\uparrow$ 60.28 64.34 $\uparrow$ 60.17 62.63 $\uparrow$ 59.32 63.98 $\uparrow$ 60.10 65.19 $\uparrow$ 61.34 61.37 $\uparrow$ 59.58 63.37 $\uparrow$
adult (n=100) 85.99 84.38 46.37 76.29 $\uparrow$ 47.30 75.74 $\uparrow$ 77.45 82.63 $\uparrow$ 80.88 81.69 $\uparrow$ 79.23 83.33 $\uparrow$ 73.52 83.38 $\uparrow$ 82.54 83.05 $\uparrow$ 76.60 82.28 $\uparrow$
drug (n=100) 83.18 80.31 79.29 79.75 $\uparrow$ 74.55 76.46 $\uparrow$ 74.54 78.01 $\uparrow$ 78.34 79.88 $\uparrow$ 73.19 79.72 $\uparrow$ 67.01 76.46 $\uparrow$ 76.92 77.57 $\uparrow$ 74.53 78.82 $\uparrow$
covid (n=200) 76.31 76.08 75.23 75.84 $\uparrow$ 72.62 74.99 $\uparrow$ 74.04 76.25 $\uparrow$ 75.28 76.82 normal-$\uparrow$ 67.08 75.62 $\uparrow$ 67.73 75.67 $\uparrow$ 74.82 75.27 $\uparrow$ 70.64 75.03 $\uparrow$
cutract (n=200) 75.04 74.25 73.48 75.28 $\uparrow$ 72.01 74.37 $\uparrow$ 71.99 74.93 $\uparrow$ 74.33 76.19 normal-$\uparrow$ 68.05 74.68 $\uparrow$ 71.47 75.67 $\uparrow$ 72.44 72.85 $\uparrow$ 68.55 74.43 $\uparrow$
maggic (n=200) 71.07 64.77 61.56 63.78 $\uparrow$ 59.76 62.53 $\uparrow$ 60.85 63.57 $\uparrow$ 58.23 58.85 $\uparrow$ 50.84 61.49 $\uparrow$ 57.52 63.00 $\uparrow$ 63.72 64.01 $\uparrow$ 58.92 62.08 $\uparrow$
seer (n=200) 89.99 88.13 86.96 87.72 $\uparrow$ 85.28 86.70 $\uparrow$ 85.05 87.83 $\uparrow$ 87.14 88.80 normal-$\uparrow$ 84.86 88.14 $\uparrow$ 80.80 87.46 $\uparrow$ 86.57 87.11 $\uparrow$ 82.30 87.59 $\uparrow$
compas (n=200) 66.22 63.61 67.68 68.80 normal-$\uparrow$ 60.67 63.43 $\uparrow$ 61.20 65.11 $\uparrow$ 60.88 63.87 $\uparrow$ 60.10 64.91 $\uparrow$ 56.94 64.76 $\uparrow$ 61.57 62.76 $\uparrow$ 59.95 64.14 $\uparrow$
adult (n=200) 86.01 85.16 40.18 77.61 $\uparrow$ 46.74 78.71 $\uparrow$ 80.92 84.26 $\uparrow$ 84.50 85.44 normal-$\uparrow$ 82.73 85.05 $\uparrow$ 77.48 84.72 $\uparrow$ 84.26 84.33 $\uparrow$ 78.12 83.94 $\uparrow$
drug (n=200) 82.14 81.47 78.70 78.98 $\uparrow$ 75.87 77.59 $\uparrow$ 77.19 80.03 $\uparrow$ 73.25 75.75 $\uparrow$ 78.50 81.13 $\uparrow$ 67.83 78.81 $\uparrow$ 79.50 80.16 $\uparrow$ 74.81 79.45 $\uparrow$
@@
cutract (n=20) 73.14 71.04 73.26 74.30 normal-$\uparrow$ 71.62 73.04 $\uparrow$ 65.13 69.67 $\uparrow$ 68.46 68.87 $\uparrow$ 51.99 68.31 $\uparrow$ 62.93 70.87 $\uparrow$ 68.28 70.72 $\uparrow$ 69.96 71.98 $\uparrow$
maggic (n=20) 71.38 57.98 61.54 62.81 normal-$\uparrow$ 58.76 60.01 $\uparrow$ 53.62 55.44 $\uparrow$ 55.97 57.06 $\uparrow$ 50.14 56.96 $\uparrow$ 54.91 58.22 $\uparrow$ 56.79 57.15 $\uparrow$ 55.41 57.72 $\uparrow$
@@
adult (n=20) 88.64 79.21 52.12 74.28 $\uparrow$ 47.70 72.91 $\uparrow$ 71.70 78.26 $\uparrow$ 76.50 77.22 $\uparrow$ 67.79 78.67 $\uparrow$ 66.79 $7$uparrow 77.11 77.97 $\uparrow$ 75.46 79.35 normal-$\uparrow$

### v3

@@ Appendix C Additional Results
C.1 Decoupling prior knowledge and data model
In-context samples Downstream accuracy
$n=1$ (Prior) $70.20\pm 1.60$
$n=20$ $73.87\pm 0.50$
$n=40$ $73.95\pm 0.67$
$n=100$ $74.71\pm 0.34$
$D_{\mathrm{oracle}}$ $74.6\pm 0.15$
From these results, we conclude that the LLM prior permits to obtain good downstream performance, but is outperformed by $\mathcal{D_{\mathrm{oracle}}}$ by a margin of $4.4\%$. Hence, we cannot solely rely on the prior. Furthermore, downstream performance increases as the number of in-context samples increases. This shows it is important to include in-context samples if we wish to obtain downstream performance close to $\mathcal{D}_{\mathrm{oracle}}$, as the LLM can build a good data model. This implies that while the LLM uses background knowledge of similar datasets, it still needs in-context samples to refine its prior by creating a good data model.
We then quantify and visualize the strength of the prior, by studying how much the LLMs output distribution adapts to the in-context samples provided. We evaluate data generated by the prior of the LLM ($n=1$), and for $n=20,40,100$ on the Covid dataset. In particular, we observe in Figure 8 that there is a region in the oracle data which is not captured by the LLM’s prior output (the left part of the leftmost blob, circled in blue in Figure 8). However, as the number of in-context real examples increases in the prompt of the LLM, we observe that this steers the LLM to generate data which covers this region. This region is associated to the subgroup of people older than $87$ years old, and having many severe comorbidities (e.g. Diabetes, Cardiovascular diseases) and many respiratory symptoms. This subgroup, in the Oracle dataset, represents less than $3.5\%$ of the data, and is completely ignored by the GPT-4 prior. In particular, the prior defaults to more typical patients in the range $70$-$80$ years old. On the contrary, as $n$ increases, the LLM is guided by the in-context samples and generates samples from this subgroup, which are ”rarer” or different from the general population.
This shows that the LLM captures the distinct features of this particular region, and is not overwhelmed by the prior. Instead, the data in the form of in-context samples adapts it, and aligns the augmented dataset with the ground-truth distribution.
C.2 Curation mechanism also improves open-source LLM generated data
In this subsection, we demonstrate the wide applicability of the CLLM  framework. Different factors may affect the choice of the LLM backbone, such as operational costs, and the desired parameterization of the LLM (for quality of generation). In addition to GPT-4 and GPT-3.5, we use the open source models Mistral-7b (Jiang et al., 2023) and LLaMa-13b (4-bit quantized) (Touvron et al., 2023), LLaMa-70b (Touvron et al., 2023) and Mixtral-8x7b (Jiang et al., 2024) to generate augmented datasets  [^3] For LLaMa-70b and Mixtral-8x7b, we only run open-source datasets due to data sharing restrictions with the endpoints for those two models.. We then compute the downstream performance when training a model on both the uncurated and curated data. As can be seen in Table 7, downstream performance with uncurated is lower for these open-source models compared to GPT-4 — which is expected given their significantly smaller size (i.e. parameter count). However, the curation mechanism, which is the key contribution of CLLM, almost always improves downstream performance for all LLM backbones investigated. Overall, this demonstrates the versatility and wide applicability of CLLM  for tabular data augmentation in low-data regimes.
Practical tip: The size of the LLM plays a role: the larger models (with more parameters) outperform those with fewer parameters (e.g. GPT-4 vs Mixtral-8x7b vs Mistral-7b). Hence, while curation helps improve all LLM-generated data, ideally the best LLM possible should be used.
Table 7: AUC averaged over 4 downstream models on $D_{\mathrm{test}}$ for GPT-4, GPT-3.5, Mistral-7b, LLAMA-13b, LLAMA-70b and Mixtral-8x7b (Curated and Uncurated)
Open AI Open source
Dataset GPT-4 GPT-3.5 Mistral-7b LLAMA-13b LLAMA-70b Mixtral
Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur.
covid (n=20) 73.78 73.87 $\uparrow$ 69.85 71.41 $\uparrow$ 71.47 72.80 $\uparrow$ 63.25 67.24 $\uparrow$ N.A. N.A. N.A. N.A.
cutract (n=20) 71.15 72.50 $\uparrow$ 69.97 71.54 $\uparrow$ 69.60 71.34 $\uparrow$ 67.84 66.71 N.A. N.A. N.A. N.A.
maggic (n=20) 60.70 61.48 $\uparrow$ 57.54 58.69 $\uparrow$ 53.64 52.06 53.30 53.96 $\uparrow$ N.A. N.A. N.A. N.A.
seer (n=20) 84.53 84.82 $\uparrow$ 83.34 83.71 $\uparrow$ 83.60 85.18 $\uparrow$ 80.94 82.84 $\uparrow$ N.A. N.A. N.A. N.A.
compas (n=20) 68.01 67.91 62.07 64.43 $\uparrow$ 56.26 59.95 $\uparrow$ 60.08 60.34 $\uparrow$ 56.10 60.68 $\uparrow$ 54.34 61.66 $\uparrow$
adult (n=20) 50.39 71.48 $\uparrow$ 49.23 72.37 $\uparrow$ 47.68 65.84 $\uparrow$ 48.82 66.00 $\uparrow$ 51.96 68.40 $\uparrow$ 57.78 76.84 $\uparrow$
drug (n=20) 75.08 75.29 $\uparrow$ 71.68 72.14 $\uparrow$ 74.14 75.21 $\uparrow$ 67.96 67.87 66.77 69.12 $\uparrow$ 72.45 71.63
covid (n=40) 73.40 73.95 $\uparrow$ 70.42 71.93 $\uparrow$ 69.61 71.47 $\uparrow$ 61.32 65.57 $\uparrow$ N.A. N.A. N.A. N.A.
cutract (n=40) 69.87 71.72 $\uparrow$ 68.47 69.56 $\uparrow$ 68.74 72.36 $\uparrow$ 64.34 67.46 $\uparrow$ N.A. N.A. N.A. N.A.
maggic (n=40) 59.29 60.77 $\uparrow$ 57.50 59.15 $\uparrow$ 52.43 53.79 $\uparrow$ 52.61 53.45 $\uparrow$ N.A. N.A. N.A. N.A.
seer (n=40) 84.29 84.93 $\uparrow$ 83.46 84.44 $\uparrow$ 83.88 85.21 $\uparrow$ 78.82 82.65 $\uparrow$ N.A. N.A. N.A. N.A.
compas (n=40) 67.57 67.85 $\uparrow$ 61.34 62.84 $\uparrow$ 57.07 60.87 $\uparrow$ 59.52 61.46 $\uparrow$ 57.48 61.20$\uparrow$ 58.90 63.32$\uparrow$
adult (n=40) 48.31 73.82 $\uparrow$ 49.21 74.27 $\uparrow$ 48.80 74.13 $\uparrow$ 54.34 69.31 $\uparrow$ 64.44 74.83$\uparrow$ 56.59 78.30 $\uparrow$
drug (n=40) 74.30 75.79 $\uparrow$ 71.33 72.76 $\uparrow$ 73.12 74.12 $\uparrow$ 69.84 72.50 $\uparrow$ 64.14 68.14 $\uparrow$ 74.34 76.07 $\uparrow$
covid (n=100) 73.77 74.71 $\uparrow$ 70.71 72.76 $\uparrow$ 71.02 73.37 $\uparrow$ 63.76 70.68 $\uparrow$ N.A. N.A. N.A. N.A.
cutract (n=100) 70.20 72.51 $\uparrow$ 69.97 71.94 $\uparrow$ 68.92 71.17 $\uparrow$ 64.81 69.85 $\uparrow$ N.A. N.A. N.A. N.A.
maggic (n=100) 58.98 61.32 $\uparrow$ 55.71 58.90 $\uparrow$ 52.53 53.36 $\uparrow$ 54.27 53.65 N.A. N.A. N.A. N.A.
seer (n=100) 84.45 85.37 $\uparrow$ 83.92 85.08 $\uparrow$ 82.23 84.36 $\uparrow$ 81.00 81.99 $\uparrow$ N.A. N.A. N.A. N.A.
compas (n=100) 68.02 68.19 $\uparrow$ 60.10 62.47 $\uparrow$ 53.74 61.28 $\uparrow$ 57.90 61.79 $\uparrow$ 60.96 63.36 $\uparrow$ 59.07 63.13 $\uparrow$
adult (n=100) 46.09 74.57 $\uparrow$ 47.56 73.97 $\uparrow$ 40.51 71.08 $\uparrow$ 48.85 72.91 $\uparrow$ 54.45 76.67 $\uparrow$ 54.27 77.61 $\uparrow$
drug (n=100) 76.24 76.74 $\uparrow$ 69.46 71.05 $\uparrow$ 74.02 76.55 $\uparrow$ 66.86 75.31 $\uparrow$ 71.47 75.00 $\uparrow$ 73.22 75.83 $\uparrow$
covid (n=200) 73.40 74.62 $\uparrow$ 70.70 73.12 $\uparrow$ 70.81 73.26 $\uparrow$ 62.53 70.67 $\uparrow$ N.A. N.A. N.A. N.A.
cutract (n=200) 71.39 73.01 $\uparrow$ 70.28 72.39 $\uparrow$ 67.69 70.29 $\uparrow$ 65.77 69.11 $\uparrow$ N.A. N.A. N.A. N.A.
maggic (n=200) 58.92 61.41 $\uparrow$ 57.33 60.16 $\uparrow$ 52.78 52.40 52.56 52.90 $\uparrow$ N.A. N.A. N.A. N.A.
seer (n=200) 84.39 85.56 $\uparrow$ 83.48 84.80 $\uparrow$ 83.15 84.18 $\uparrow$ 80.88 82.74 $\uparrow$ N.A. N.A. N.A. N.A.
compas (n=200) 67.02 68.15 $\uparrow$ 60.48 63.39 $\uparrow$ 53.96 59.22 $\uparrow$ 57.56 59.97 $\uparrow$ 61.82 63.53 $\uparrow$ 60.99 64.72 $\uparrow$
adult (n=200) 40.96 75.84 $\uparrow$ 49.89 76.11 $\uparrow$ 44.13 74.23 $\uparrow$ 48.25 78.35 $\uparrow$ 54.42 76.23 $\uparrow$ 42.42 76.06 $\uparrow$
drug (n=200) 75.58 76.06 $\uparrow$ 70.66 72.81 $\uparrow$ 71.89 76.54 $\uparrow$ 70.88 75.46 $\uparrow$ 69.12 74.22 $\uparrow$ 73.43 76.31 $\uparrow$
C.3 Ablation for contextual information on Compas
We conduct a similar experiment as in Table 2, and use the dataset Compas. We report the results in Table C.3.
P R U P R U P R U
20 $\bf 0.69_{\scriptsize(0.02)}$ $\bf 0.88_{\scriptsize(0.02)}$ $\bf 0.69_{\scriptsize(0.02)}$ $0.27_{\scriptsize(0.03)}$ $\bf 0.89_{\scriptsize(0.03)}$ $0.60_{\scriptsize(0.03)}$ $0.43_{\scriptsize(0.02)}$ $0.43_{\scriptsize(0.05)}$ $0.55_{\scriptsize(0.04)}$
40 $\bf 0.70_{\scriptsize(0.0)}$ $\bf 0.92_{\scriptsize(0.01)}$ $\bf 0.65_{\scriptsize(0.03)}$ $0.31_{\scriptsize(0.06)}$ $0.84_{\scriptsize(0.03)}$ $0.57_{\scriptsize(0.01)}$ $0.54_{\scriptsize(0.02)}$ $0.80_{\scriptsize(0.02)}$ $0.50_{\scriptsize(0.04)}$
100 $\bf 0.69_{\scriptsize(0.02)}$ $\bf 0.89_{\scriptsize(0.02)}$ $\bf 0.69_{\scriptsize(0.01)}$ $0.34_{\scriptsize(0.1)}$ $0.85_{\scriptsize(0.05)}$ $0.62_{\scriptsize(0.01)}$ $0.60_{\scriptsize(0.03)}$ $0.86_{\scriptsize(0.02)}$ $0.59_{\scriptsize(0.03)}$
200 $\bf 0.70_{\scriptsize(0.01)}$ $\bf 0.89_{\scriptsize(0.02)}$ $\bf 0.69_{\scriptsize(0.01)}$ $0.31_{\scriptsize(0.05)}$ $0.87_{\scriptsize(0.03)}$ $0.58_{\scriptsize(0.05)}$ $0.65_{\scriptsize(0.02)}$ $0.88_{\scriptsize(0.01)}$ $0.63_{\scriptsize(0.01)}$
C.4 Comparison to random noise baseline
C.5 Detailed results for Section 3.1
@@ C.4 Comparison to random noise baseline
We report additional results for Sec. 3.1, showing the AUC for each downstream model (XGBoost, Random forest, Logistic regression, Decision tree). As we can see, the conclusion that curation helps improve downstream performance holds for each of these various downstream models, as is indicated by the green arrows in Tables 9, 10, 11, 12.
@@
cutract (n=20) 74.16 73.34 73.45 74.24 $\uparrow$ 72.22 73.51 $\uparrow$ 66.24 70.45 $\uparrow$ 69.73 70.49 $\uparrow$ 52.61 68.11 $\uparrow$ 66.63 72.39 $\uparrow$ 71.42 72.34 $\uparrow$ 71.18 71.89 $\uparrow$
maggic (n=20) 71.28 59.18 62.92 63.97 $\uparrow$ 60.10 61.25 $\uparrow$ 53.21 55.14 $\uparrow$ 55.61 56.50 $\uparrow$ 50.92 57.23 $\uparrow$ 55.91 59.07 $\uparrow$ 58.20 58.55 $\uparrow$ 55.43 57.71 $\uparrow$
seer (n=20) 90.09 85.22 86.30 86.82 $\uparrow$ 85.86 85.73 78.11 80.52 $\uparrow$ 82.61 82.74 $\uparrow$ 47.57 75.70 $\uparrow$ 77.35 81.45 $\uparrow$ 82.09 83.00 $\uparrow$ 77.63 81.16 $\uparrow$
@@
covid (n=40) 76.11 75.21 75.63 75.70 $\uparrow$ 72.27 73.59 $\uparrow$ 66.64 70.75 $\uparrow$ 75.56 75.39 56.91 70.88 $\uparrow$ 70.47 73.15 $\uparrow$ 73.59 73.63 $\uparrow$ 64.27 69.71 $\uparrow$
cutract (n=40) 74.16 72.31 71.67 73.20 $\uparrow$ 69.92 71.01 $\uparrow$ 64.67 69.21 $\uparrow$ 69.16 69.67 $\uparrow$ 53.41 68.93 $\uparrow$ 61.23 69.21 $\uparrow$ 70.82 71.21 $\uparrow$ 59.99 67.56 $\uparrow$
maggic (n=40) 71.28 60.91 61.80 63.11 $\uparrow$ 59.38 61.50 $\uparrow$ 56.49 58.17 $\uparrow$ 56.50 58.21 $\uparrow$ 48.43 57.84 $\uparrow$ 55.47 59.94 $\uparrow$ 59.82 60.11 $\uparrow$ 56.76 58.92 $\uparrow$
seer (n=40) 90.09 86.96 86.01 86.29 $\uparrow$ 86.73 87.09 $\uparrow$ 83.46 86.62 $\uparrow$ 85.86 85.19 54.71 83.22 $\uparrow$ 83.23 86.81 $\uparrow$ 83.76 84.82 $\uparrow$ 80.34 85.82 $\uparrow$
compas (n=40) 66.79 62.73 68.06 68.13 $\uparrow$ 61.17 62.27 $\uparrow$ 56.05 60.92 $\uparrow$ 59.76 61.03 $\uparrow$ 57.82 64.25 $\uparrow$ 58.89 63.79 $\uparrow$ 61.20 61.25 $\uparrow$ 55.89 60.54 $\uparrow$
adult (n=40) 85.83 83.61 50.26 75.55 $\uparrow$ 46.68 76.63 $\uparrow$ 75.13 81.56 $\uparrow$ 76.10 80.40 $\uparrow$ 68.04 81.08 $\uparrow$ 73.29 81.46 $\uparrow$ 80.35 81.10 $\uparrow$ 75.81 82.11 $\uparrow$
drug (n=40) 83.12 78.81 78.35 79.54 $\uparrow$ 77.33 77.94 $\uparrow$ 74.68 76.64 $\uparrow$ 74.50 77.18 $\uparrow$ 70.25 $7$uparrow 65.86 74.55 $\uparrow$ 74.84 75.26 $\uparrow$ 71.29 75.98 $\uparrow$
covid (n=100) 76.11 75.78 75.86 76.33 $\uparrow$ 73.02 74.40 $\uparrow$ 72.00 74.74 $\uparrow$ 74.64 75.74 $\uparrow$ 65.76 74.69 $\uparrow$ 67.05 75.47 $\uparrow$ 74.00 74.09 $\uparrow$ 72.11 74.44 $\uparrow$
cutract (n=100) 74.16 73.93 72.92 74.73 $\uparrow$ 72.51 73.93 $\uparrow$ 70.18 74.26 $\uparrow$ 71.71 73.69 $\uparrow$ 55.48 70.79 $\uparrow$ 69.79 74.75 $\uparrow$ 71.18 72.47 $\uparrow$ 68.47 73.78 $\uparrow$
maggic (n=100) 71.28 63.06 60.99 63.36 $\uparrow$ 57.97 60.86 $\uparrow$ 58.98 60.76 $\uparrow$ 58.88 60.13 $\uparrow$ 49.66 59.57 $\uparrow$ 57.82 62.24 $\uparrow$ 61.55 61.90 $\uparrow$ 57.46 60.55 $\uparrow$
seer (n=100) 90.00 87.53 86.33 87.31 $\uparrow$ 86.40 87.06 $\uparrow$ 84.59 87.27 $\uparrow$ 85.89 87.00 $\uparrow$ 70.32 85.97 $\uparrow$ 84.30 87.52 $\uparrow$ 85.12 85.82 $\uparrow$ 81.92 86.41 $\uparrow$
compas (n=100) 66.79 63.18 68.44 68.67 $\uparrow$ 59.41 62.00 $\uparrow$ 60.28 64.34 $\uparrow$ 60.17 62.63 $\uparrow$ 59.32 63.98 $\uparrow$ 60.10 65.19 $\uparrow$ 61.34 61.37 $\uparrow$ 59.58 63.37 $\uparrow$
adult (n=100) 85.83 84.38 46.37 76.29 $\uparrow$ 47.30 75.74 $\uparrow$ 77.45 82.63 $\uparrow$ 80.88 81.69 $\uparrow$ 79.23 83.33 $\uparrow$ 73.52 83.38 $\uparrow$ 82.54 83.05 $\uparrow$ 76.60 82.28 $\uparrow$
drug (n=100) 83.12 80.31 79.29 79.75 $\uparrow$ 74.55 76.46 $\uparrow$ 74.54 78.01 $\uparrow$ 78.34 79.88 $\uparrow$ 73.19 79.72 $\uparrow$ 67.01 76.46 $\uparrow$ 76.92 77.57 $\uparrow$ 74.53 78.82 $\uparrow$
covid (n=200) 76.11 76.08 75.23 75.84 $\uparrow$ 72.62 74.99 $\uparrow$ 74.04 76.25 $\uparrow$ 75.28 76.82 $\uparrow$ 67.08 75.62 $\uparrow$ 67.73 75.67 $\uparrow$ 74.82 75.27 $\uparrow$ 70.64 75.03 $\uparrow$
cutract (n=200) 74.16 74.25 73.48 75.28 $\uparrow$ 72.01 74.37 $\uparrow$ 71.99 74.93 $\uparrow$ 74.33 76.19 $\uparrow$ 68.05 74.68 $\uparrow$ 71.47 75.67 $\uparrow$ 72.44 72.85 $\uparrow$ 68.55 74.43 $\uparrow$
maggic (n=200) 71.28 64.77 61.56 63.78 $\uparrow$ 59.76 62.53 $\uparrow$ 60.85 63.57 $\uparrow$ 58.23 58.85 $\uparrow$ 50.84 61.49 $\uparrow$ 57.52 63.00 $\uparrow$ 63.72 64.01 $\uparrow$ 58.92 62.08 $\uparrow$
seer (n=200) 90.09 88.13 86.96 87.72 $\uparrow$ 85.28 86.70 $\uparrow$ 85.05 87.83 $\uparrow$ 87.14 88.80 $\uparrow$ 84.86 88.14 $\uparrow$ 80.80 87.46 $\uparrow$ 86.57 87.11 $\uparrow$ 82.30 87.59 $\uparrow$
compas (n=200) 66.79 63.61 67.68 68.80 $\uparrow$ 60.67 63.43 $\uparrow$ 61.20 65.11 $\uparrow$ 60.88 63.87 $\uparrow$ 60.10 64.91 $\uparrow$ 56.94 64.76 $\uparrow$ 61.57 62.76 $\uparrow$ 59.95 64.14 $\uparrow$
adult (n=200) 85.83 85.16 40.18 77.61 $\uparrow$ 46.74 78.71 $\uparrow$ 80.92 84.26 $\uparrow$ 84.50 85.44 $\uparrow$ 82.73 85.05 $\uparrow$ 77.48 84.72 $\uparrow$ 84.26 84.33 $\uparrow$ 78.12 83.94 $\uparrow$
drug (n=200) 83.12 81.47 78.70 78.98 $\uparrow$ 75.87 77.59 $\uparrow$ 77.19 80.03 $\uparrow$ 73.25 75.75 $\uparrow$ 78.50 81.13 $\uparrow$ 67.83 78.81 $\uparrow$ 79.50 80.16 $\uparrow$ 74.81 79.45 $\uparrow$
@@
cutract (n=20) 73.14 71.04 73.26 74.30 $\uparrow$ 71.62 73.04 $\uparrow$ 65.13 69.67 $\uparrow$ 68.46 68.87 $\uparrow$ 51.99 68.31 $\uparrow$ 62.93 70.87 $\uparrow$ 68.28 70.72 $\uparrow$ 69.96 71.98 $\uparrow$
maggic (n=20) 71.38 57.98 61.54 62.81 $\uparrow$ 58.76 60.01 $\uparrow$ 53.62 55.44 $\uparrow$ 55.97 57.06 $\uparrow$ 50.14 56.96 $\uparrow$ 54.91 58.22 $\uparrow$ 56.79 57.15 $\uparrow$ 55.41 57.72 $\uparrow$
@@
adult (n=20) 88.64 79.21 52.12 74.28 $\uparrow$ 47.70 72.91 $\uparrow$ 71.70 78.26 $\uparrow$ 76.50 77.22 $\uparrow$ 67.79 78.67 $\uparrow$ 66.79 $7$uparrow 77.11 77.97 $\uparrow$ 75.46 79.35 $\uparrow$